Suppose that speeding fines are determined by the linear function
​f(x)= 10 (x-65)+50,xgreat then65, where​ f(x) is the cost in dollars of the fine if a person is caught driving x miles per hour.
​(a) Radar clocked a driver at 76 mph. How much was the​ fine?
​(b) While balancing his​ checkbook, a person ran across a canceled check that his wife had written to the Department of Motor Vehicles for a speeding fine. The check was written for​ $100. How fast was his wife​ driving?
​(c) At what​ whole-number speed are tickets first​ given?
​(d) For what speeds is the fine greater than​ $200?
(a) The fine was $

Answer by stanbon(75887)

Suppose that speeding fines are determined by the linear function
f(x)= 10 (x-65)+50,(x greater then 65), where f(x) is the cost in dollars of the fine if a person is caught driving x miles per hour.
(a) Radar clocked a driver at 76 mph. How much was the fine?
f(76) = 10(76-65)+50 = $160.00
-----------------------------
​(b) While balancing his checkbook, a person ran across a canceled check that his wife had written to the Department of Motor Vehicles for a speeding fine. The check was written for $100. How fast was his wife driving?
100 = 10(x-65)+50
50 = 10(x-65)
5 = x-65
x = 70 mph
--------------
(c) At what whole-number speed are tickets first given?
10(x-65)+50 >= 0
x-65 > - 5
x >= 70
-------
(d) For what speeds is the fine greater than $200?
10(x-65)+50 > 200
x-65 > 15
x > 80 mph
--------------
Cheers,
Stan H.
--------